Please help on this problem. solve for x.
2x + 3y=.78
3x + 2y=.72
(In a stationery store, pencils have one price and pens have another. Two pencils and three pens cost 78¢. But three pencils and two pens cost 72¢. How much does one pencil cost?)
Thanks!!!

Answer by stanbon(75887)

2x + 3y=.78
3x + 2y=.72
(In a stationery store, pencils have one price and pens have another. Two pencils and three pens cost 78¢. But three pencils and two pens cost 72¢. How much does one pencil cost?)
1st: 2x + 3y=.78
2nd: 3x + 2y=.72
---------------------
Multiply 1st by 3 and 2nd by 2 to get:
3rd: 6x + 9y = 2.34
4th: 6x + 4y = 1.44
---------------------
Subtract 4th from 3rd to solve for y:
5y = 0.90
y = 0.18 (price of a pen)
--------------
Substitute that into 1st (or 2nd) to solve for "x":
2x + 3*0.18 = 78
2x + 0.54 = 0.78
2x = 0.24
x = 0.12 (price of a pencil)
===========
Cheers,
Stan H.